OpenStudy (anonymous):

Write a quadratic equation in STANDARD FORM with given solution set: {1/6, 2/3}

OpenStudy (anonymous):

(6x - 1)(3x - 2) = 0 18x^2 - 12x - 3x + 2 = 0 18x^2 - 15x + 2 = 0

OpenStudy (anonymous):

So what are the possible answers then?